DETECTION OF CHANGES IN LINEAR SEQUENCES

LAJOS HORVÁTH

Department of Mathematics, University of Utah, Salt Lake City, UT 84112, U.S.A.

(Received May 15, 1995; revised May 1, 1996)

Abstract. We discuss the asymptotic properties of some tests to detect possible changes in the mean of linear processes.

Key words and phrases: Change-point, linear sequence, dependent observations.